how much gas money to drive 211 miles
Found 2 solutions by JulietG, richwmiller:
Answer by JulietG(1812)

It depends upon how many miles per gallon your car gets. Typically, in city driving, about 20 mpg.
211 / 20 = 10.55 gallons. At today's rate of ~$4 per gallon = $42.20
.
If you happen to have a smaller car or are driving on the freeway, let's say you get closer to 27 miles per gallon. 211/27=7.81 gallons. Maybe you don't have the high taxes I do and your gas is "only" $3.59 gallon. Multiply # of gallons by price.
7.81*3.59 = $28.04

Answer by richwmiller(17219)

Zero. I drive an electric car.